OpenStudy (ashleyisakitty):

Given the proportion a/b=7/13, what ratio completes the equivalent proportion a/7=? I think it is b/13.

OpenStudy (***[isuru]***):

yes u r correct !!!

OpenStudy (nincompoop):

correct

OpenStudy (jdoe0001):

sounds good, yes

OpenStudy (ashleyisakitty):

Thanks all :)

OpenStudy (***[isuru]***):

u r welcome

OpenStudy (luigi0210):

Can I be the idiot to say it's wrong?

OpenStudy (nincompoop):

\[\frac{ a }{ b }=\frac{ 7 }{ 13 }\] \[\frac{ a }{ 7 }=1=\frac{ b }{ 13 }\]